physics help... detailed solution pls


A black thin-walled cylindrical can, 10.0 cm in di
Answer #1

Twater=100 deg celcius = 373 K

Troom= 20 deg celcius = 293 K.

Area=2pi*r(r+h)=0.07854 m^2

Volume(V) of cylinder=pi*r^2*h=15.7*10^-4 m^3

mass of water=pV=1.57 kg


m*C*dT/dt=AS(Twater^4 - Troom^4) where S is boltzman constant=5.67*10^-8 Wm^2/K^4

1.57*4186*dT/dt=0.07854*5.67*10^-8(373^4-293^4)


So,dT/dt=0.008122 K/s.
Answer #2

A black thin-walled cylindrical can, 10.0 cm in diameter and 20.0 cm tall is filled with water at 100